OpenStudy (anonymous):

h -2 and 3/4=1and 1/2 so what does h=

OpenStudy (anonymous):

helppppp

jimthompson5910 (jim_thompson5910):

the problem is this right? \[\Large h - 2\frac{3}{4} = 1\frac{1}{2}\]

OpenStudy (anonymous):

I just want to know what h =

jimthompson5910 (jim_thompson5910):

I know, but I need to know what the actual problem is

OpenStudy (anonymous):

what do you mean?

jimthompson5910 (jim_thompson5910):

is the actual problem \[\Large h - 2\frac{3}{4} = 1\frac{1}{2}\]

OpenStudy (anonymous):

yes that is the actual problem

jimthompson5910 (jim_thompson5910):

ok you would add 2 3/4 to both sides to get this \[\Large h - 2\frac{3}{4} = 1\frac{1}{2}\] \[\Large h = 1\frac{1}{2} + 2\frac{3}{4}\] \[\Large h = ??\]

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):